# Quantitative Ultrasound for Periodontal Soft Tissue Characterization

### v1

@@ Conclusions
This study is among early investigations in the literature into applications of QUS approaches for periodontal soft tissue characterizations and serves as a crucial preliminary step with promising results towards employing QUS as an additional diagnostic tool for disease assessment in periodontics. In this study, characterization of periodontal soft tissues (alveolar mucosa and gingiva) was investigated in an in vivo swine model using a QUS approach based on the Burr and Nakagami models for speckle statistics. The results showed that the Burr parameters (power-law parameter $b$ and the scale factor $l$) and also Nakagami parameters (shape parameter $m$ and scale factor $\alpha$) have potentials to distinguish clinically significant tissue types. This study demonstrated that the Burr power-law parameter and Nakagami shape parameter were significantly higher in gingiva compared to alveolar mucosa while the Burr scale factor and the Nakagami scale factor were significantly lower in gingiva. The QUS findings were hypothesized to be aligned with qualitative assessments of histology using Masson’s Trichrome and H&E staining techniques. The two tissue types were classified in 2D parameter spaces using the QUS parameters from the Burr and Nakagami models which yielded a separation accuracy of 93.51% and 90.91%, respectively. The classification of the two tissue types using parameters from the two models in 4D resulted in a classification accuracy of 92.21%. Further studies should assess the effect of disease conditions such as oral soft tissue inflammation on QUS parameters. Our results indicate that QUS could potentially become an augmented tool in periodontics, as an objective, quantitative and noninvasive technique for disease diagnosis, longitudinal monitoring of healing, and feedback for indicated interventions.
@@
Sedigheh Poul: Conceptualization, Methodology, Histology Imaging, Formal Analyses and Software, Validations and Visualization, Writing Original Draft and Revision;
@@
Hsun-Liang Chan, Oliver D. Kripfgans: Conceptualization, Methodology, Animal Experiment, US Scan Data Acquisition, Funding Acquisition, Review and Editing;

### v2

@@ Conclusions
This study is among early investigations in the literature into applications of QUS approaches for periodontal soft tissue characterizations and serves as a crucial preliminary step with promising results towards employing QUS as an additional diagnostic tool for disease assessment in periodontics. In this study, characterization of periodontal soft tissues (alveolar mucosa and gingiva) was investigated in an in vivo swine model using a QUS approach based on the Burr and Nakagami models for speckle statistics. The results showed that the Burr parameters (power-law parameter $b$ and the scale factor $l$) and also Nakagami parameters (shape parameter $m$ and scale factor $\Omega$) have potentials to distinguish clinically significant tissue types. This study demonstrated that the Burr power-law parameter and Nakagami shape parameter were significantly higher in gingiva compared to alveolar mucosa while the Burr scale factor and the Nakagami scale factor were significantly lower in gingiva. The QUS findings were hypothesized to be aligned with qualitative assessments of histology using Masson’s Trichrome and H&E staining techniques. The two tissue types were classified in 2D parameter spaces using the QUS parameters from the Burr and Nakagami models which yielded a separation accuracy of 93.51% and 90.91%, respectively. The classification of the two tissue types using parameters from the two models in 4D resulted in a classification accuracy of 92.21%. Further studies should assess the effect of disease conditions such as oral soft tissue inflammation on QUS parameters. Our results indicate that QUS could potentially become an augmented tool in periodontics, as an objective, quantitative and noninvasive technique for disease diagnosis, longitudinal monitoring of healing, and feedback for indicated interventions.
@@
Daria Poul: Conceptualization, Methodology, Histology Imaging, Formal Analyses and Software, Validations and Visualization, Writing Original Draft and Revision;
@@
Hsun-Liang Chan, Oliver D. Kripfgans: Conceptualization, Methodology, Animal Experiment, US Scan Data Acquisition, Funding Acquisition, Review and Editing.
Conflicts of Interest
All authors certify that they have no conflicts of interest to report.
@@ Acknowledgements
Data Availability Statement
Inquiries regarding the availability of data supporting the findings of this study should be directed to the corresponding authors.